OpenStudy (anonymous):

What is the axis of symmetry of the parabola represented by the function y=2x2 + 12x + 10?

OpenStudy (raffle_snaffle):

Graph it.

OpenStudy (raffle_snaffle):

It's going to be a parabola.

OpenStudy (raffle_snaffle):

Here this may help=) http://www.mathwarehouse.com/geometry/parabola/axis-of-symmetry.php

OpenStudy (anonymous):

I dont know how to graph equations

OpenStudy (raffle_snaffle):

You can do it one of two ways. You can either graph it by using a graphing calculator or you can make a table.

OpenStudy (anonymous):

i dont know how to do it like that, my calculator doesnt have an x, does it?

OpenStudy (campbell_st):

just use for a parabola \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\]

OpenStudy (campbell_st):

in your question b = 12 and a = 2

OpenStudy (campbell_st):

a 3rd method is to factor the equation and the line of symmetry is half way between the zeros

OpenStudy (raffle_snaffle):

Campbell_st how did you obtain the equation x=-b/2a did you solve for x?

OpenStudy (anonymous):

Oh my gosh, im so sorry but i'm really confused :/

OpenStudy (campbell_st):

its just a know fact.... the general quadratic formula gives the zeros... and x = -b/2a is the midpoint between the zeros... I learnt it is year 10

OpenStudy (raffle_snaffle):

It has been quite awhile for me as well.. The formula looks familiar from a previous class...

OpenStudy (campbell_st):

ok... so use the equation for any parabola \[y =ax^2 + bx + c\] the line of symmetry can be found using \[x = \frac{-b}{2a}\] so in your question, a = 2 and b = 12 so its \[x = \frac{-12}{2 \times 2}\] so the line of symmetry is x = -3... which makes sense as the zeros are x = -5 and x = -2

OpenStudy (isaiah.feynman):

Mind if I butt in @AmberLynn127 The axis of symmetry is the y axis, because the function is even. Even functions are symmetric about the y axis.

OpenStudy (campbell_st):

and the reason it looks familiar is that the general quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] take out the radical and you have \[x = \frac{-b}{2a}\]

OpenStudy (anonymous):

i know that! i just dont know how to make the graph, if i had the graph i could figure it out

OpenStudy (raffle_snaffle):

So @campbell_st do you obtain x=-b/2a by deriving the equation y=ax^3+bx+c

OpenStudy (campbell_st):

the function isn't even f(-x) = 2(-x)^2 + 12(-x) + 10 which is f(x) = 2x^2 - 12x + 10 so f(-x) does not equal f(x) the function is neither odd nor even

OpenStudy (isaiah.feynman):

@AmberLynn127 That's good for you. There is the procedue to graph quadratic functions Find x intercepts Find y intercepts Find the co ordinates of the vertex.

OpenStudy (anonymous):

hey wait!! so i have another one, and i dont remember how to set up the problem to solve it, so any help? Michelle has a garden in the shape of a square. How long is one side of the square garden if it has an area of 144 ft2?

OpenStudy (campbell_st):

well for any parabola... y= ax^2 + bx + c x = -b/2a will always give the line of symmetry then if the curve is \[x = ay^2 + by + c\] the line of symmetry is \[y = \frac{-b}{2a}\]

OpenStudy (isaiah.feynman):

@campbell_st I didn't realize that the function is not even.

OpenStudy (raffle_snaffle):

Okay thanks @campbell_st

OpenStudy (campbell_st):

if you look at it, it can't be symmetry about the y axis... is it was then the parabola would be the difference of 2 squares or y = ax^2

OpenStudy (campbell_st):

well in a square all sides are equal so then \[144 = l^2\] where l is the side length take the square root and use the positive answer.. you can't have a negative side length

OpenStudy (anonymous):

so the square root of 144?

OpenStudy (mathmale):

The axis of symmetry is the vertical line through x=-b/(2a), which in this case is not x=0 (the y-axis). x=-b/(2a) is also the coordinate of the vertex. As Campbell has correctly pointed out, this x coordinate is x=-12/(2*2) = -3. To find the vertex, substitute this x value (-3) into the original quadratic and calculate the y value. Then you'll have the coordinates of the vertex.

OpenStudy (campbell_st):

thats correct

OpenStudy (mathmale):

Amber-Lynn: mind summarizing now what you know and what you still need to figure out?

OpenStudy (campbell_st):

@mathmale thats provided its not a parabola in the form y = ay^2 + by + c

OpenStudy (anonymous):

Its just hard for me to remember every thing because school is CRAZY and want us to memorize every single equasion form in a week!

OpenStudy (mathmale):

Campbell? You don't really want y on both sides of that equation, do you? Unsure of what you're trying to say.

OpenStudy (mathmale):

Amber-L: I sympathize. But I do still suggest you read over the discussion you and several others have had about your original question and then list what you've learned and what you still need to learn to answer that question.

OpenStudy (anonymous):

i have not learned a thing. It didn't help. They didn't explain anything

OpenStudy (isaiah.feynman):

At least you know even and odd functions.

OpenStudy (anonymous):

Is everyone talking to me? or each other? cause i thought you guys were talking to each other,,

OpenStudy (isaiah.feynman):

Actually I referred to you @AmberLynn127

OpenStudy (anonymous):

IM LOST

OpenStudy (mathmale):

A-L: Sorry you feel lost. Several people have invested considerable time and thought into helping you address the problem you need to solve. I was asking that you make the effort to skim through what they have written and summarize their input for your own understanding of this problem. Were you able to do this, you'd be in a correspondingly better position to explain what it is that you still don't understand. Try to put yourself into the position of those who tried to help you, only to be told by you, "i have not learned a thing. It didn't help. They didn't explain anything." I won't be adding any more to this discussion. Good luck.

OpenStudy (anonymous):

Umm actually i read everything, but it seemed like they were on whole nother topic.

OpenStudy (mathmale):

Perhaps you need to read this material again, critically, deciding for yourself what's helpful and what's not. Others can do only so much for you; at some point you yourself have to step in, get involved, spend the time and effort, and take control.